OpenStudy (anonymous):

help me plzzz Solve for y. y/4-(y-1)/2=>3

OpenStudy (aum):

Is this the inequality? \(\Large \frac{y}{4} + \frac{y-1}{2} \ge 3 \)

OpenStudy (anonymous):

yes

OpenStudy (aum):

I realized I made a mistake in the sign. Is this the one? Check carefully \(\Large \frac{y}{4} - \frac{y-1}{2} \ge 3\)

OpenStudy (anonymous):

oh my bad it look good. :)

OpenStudy (aum):

Okay. Multiply throughout by 4 to get rid of the fraction: \(\large \frac{y}{4}*4 - \frac{y-1}{2}*4 \ge 3*4 \) = \(\large y - 2(y-1) \ge 12\) \(\large y - 2y+2 \ge 12\) \(\large -y+2 \ge 12\) \(\large -y \ge 10\) \(\large y \le 10\)

OpenStudy (aum):

The last line should be: \(\Large y \le -10\) That is, y is less than or equal to minus 10.

OpenStudy (anonymous):

hey thanks

OpenStudy (aum):

you are welcome.